OpenStudy (anonymous):

find four consecutive even integers such that -4 times the sum of the first and the fourth is six greater than the opposite of the sum of the second and the fourth

OpenStudy (anonymous):

Let the consecutive even integers be x, x+2, x+4, x+6

OpenStudy (anonymous):

ok ik how to get that far... its the solving i dont understand

OpenStudy (anonymous):

-4(x+x+6)=6[-(x+2+x+6)

OpenStudy (anonymous):

so it would be -4x+-4x+-24=6x+12+6x+36 right??

OpenStudy (anonymous):

to avoid problems solve, inside brackets first, so left hand side would be -4(2x+6) which is -8x-24

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve for x, and then plug in and find other integers, got it?

OpenStudy (anonymous):

uhmm i think so not sure

OpenStudy (anonymous):

So far you did good, except on the right hand side, you left out the negative sign

OpenStudy (anonymous):

oh wait where does the negative sign go then???

OpenStudy (anonymous):

So far you have -8x-24=-12x-48

OpenStudy (anonymous):

ok so add -8x to -12x and add -48 to -24 right??

OpenStudy (anonymous):

i got x=-6

OpenStudy (anonymous):

Yes, you ended up with the right answer, be able to show your work in your home work. Punch -6 into the four integers all the way at the top that we started with.

OpenStudy (anonymous):

so x=-6 -6+2 -6+4 and -6+6

OpenStudy (anonymous):

oh btw thank yhu(: